# Long-Context Language Models?

### v2

@@ 7 Conclusion
We present Ruler, a synthetic benchmark for evaluating long-context language models. Ruler contains diverse task categories, retrieval, multi-hop tracing, aggregation and question answering, providing a flexible and comprehensive evaluation of LLM’s long-context capabilities. We benchmark ten long-context LMs using Ruler with context sizes ranging from 4K to 128K. Despite achieving perfect results in the widely used needle-in-a-haystack test, all models fail to maintain their performance in other tasks of Ruler as we increase input length. We observe common failure modes at large context sizes, including the failure to ignore distractors and ineffective utilization of long context (e.g., simply copy from context or use parametric knowledge instead). We show that Ruler is challenging for even the top-ranked open-source models as we increase task complexity. Our analysis further reveals the large potential for improvement on Ruler and the benefit of scaling model sizes in achieving better long context capabilities.
@@ Appendix A Models
We select in total 30 models for evaluation and analysis. Our results in the main text only include aligned models (one closed-source model GPT-4 and 9 open-source models). Besides the aligned models, we also evaluate 7 open-source base models using Ruler. We use the performance of Llama2-7b (base) and Llama2-7b (chat) at context length of 4k as the threshold for determining effective context size. In our analysis section, we evaluate in total 11 models, including model series Yi and LWM, as well as models of novel architectures, including Mamba and RWKV.
@@
Command-R (Cohere, 2024) 35B 128K CohereForAI/c4ai-command-r-v01
@@
Mixtral (Jiang et al., 2024) 8x7B 32K mistralai/Mixtral-8x7B-Instruct-v0.1
Mistral (Mistral.AI, 2023) 7B 32K mistralai/Mistral-7B-Instruct-v0.2
ChatGLM (Du et al., 2022) 6B 128M THUDM/chatglm3-6b-128k
@@
Yi series (Young et al., 2024) 6B/9B 200K 01-ai/Yi-6B-200K, 01-ai/Yi-9B-200K
LWM series (Liu et al., 2024a) 7B 128/256/512K LargeWorldModel/LWM-Text-Chat-128/256/512K
LWM-base series (Liu et al., 2024a) 7B 32/128/256/512K LargeWorldModel/LWM-Text-32/128/256/512K
@@ Appendix B Task Configurations
  • Retrieval: In S-NIAH, we include the passkey retrieval (Mohtashami & Jaggi, 2023) and the vanilla NIAH (Kamradt, 2023), both use word-number as key-value and differ only by the background haystack. Additionally, we change the value type to UUID, for the purpose of testing model robustness at retrieving long strings from context. For MK-NIAH, we add three distractor needles into the haystack. We also include existing setups from previous works: line retrieval (Li et al., 2023a) and key-value retrieval (Liu et al., 2024c) with the haystack filled entirely with distractor needles. For MV-NIAH and MQ-NIAH, we test 4 values and queries respectively.
@@
type_key = word
type_value = number
type_haystack = repeat
@@
type_key = word
type_value = number
type_haystack = essay
[…]
type_key = word
type_value = uuid
type_haystack = essay
MK-NIAH
num_keys = 4
type_key = word
type_value = number
type_haystack = essay
num_keys = full haystack
type_key = word
type_value = number
[…]
num_keys = full haystack
type_key = uuid
type_value = uuid
@@
QA
dataset = SQuAD
dataset = HotpotQA
@@ Appendix C Task Correlation Analysis
Ruler is designed under the assumption that tasks across different categories are able to reveal distinct model behaviors. We conduct a preliminary correlational study to confirm the validity of task categories and guide the selection of representative tasks. We evaluate eight models (not including GPT-4 and Command-R) at various context sizes across 18 task configurations. Each task can then be represented with a vector of model performance at various context sizes. The 18 task vectors are then clustered via agglomorative clustering algorithm, using correlation coefficient as the distance metric. As shown in Figure 5, while certain tasks exhibit moderate correlations with others, tasks in each of the four categories (NIAH, VT, AG, QA) form cohesive clusters of their own without redundancy. We further eliminate a few tasks that correlate highly with other tasks within the same cluster, and finalize 13 tasks for later large scale evaluation.
@@ Appendix D Prompt Templates
GPT-4
{task_template} Do not provide any explanation. Please directly give me the answer. {task_answer_prefix}
Yi/Base
{task_template} {task_answer_prefix}
Command-R
@@
$\langle$|⟨ |START_OF_TURN_TOKEN|$\rangle$| ⟩
@@
$\langle$|⟨ |END_OF_TURN_TOKEN|$\rangle$| ⟩
$\langle$|⟨ |START_OF_TURN_TOKEN|$\rangle$| ⟩
@@
LWM
You are a helpful assistant. USER: {task_template} ASSISTANT: {task_answer_prefix}
LongChat
A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. USER: {task_template} ASSISTANT: {task_answer_prefix}
ChatGLM
[ gMASK] sop|$\langle$|⟨ |user|$\rangle$| ⟩
@@
{task_answer_prefix}
Others
[ INST]  {task_template} [ /INST]  {task_answer_prefix}
@@
Task Template:
Some special magic numbers are hidden within the following text. Make sure to memorize it. I will quiz you about the numbers afterwards.
The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again.
…… One of the special magic numbers for {word} is: {number}. ……
What is the special magic number for {word} mentioned in the provided text?
Task Answer Prefix:
The special magic number for {word} mentioned in the provided text is
@@
Task Template:
Some special magic numbers are hidden within the following text. Make sure to memorize it. I will quiz you about the numbers afterwards.
Paul Graham Essays.
…… One of the special magic numbers for {word} is: {number}. ……
What is the special magic number for {word} mentioned in the provided text?
Task Answer Prefix:
The special magic number for {word} mentioned in the provided text is
@@
Task Template:
Some special magic words are hidden within the following text. Make sure to memorize it. I will quiz you about the words afterwards.
Paul Graham Essays.
…… One of the special magic words for {word} is: {word}. ……
What is the special magic word for {word} mentioned in the provided text?
Task Answer Prefix:
The special magic word for {word} mentioned in the provided text is
@@
Task Template:
Some special magic numbers are hidden within the following text. Make sure to memorize it. I will quiz you about the numbers afterwards.
Paul Graham Essays.
…… One of the special magic numbers for {word-1} is: {number-1}. ……
…… One of the special magic numbers for {word-2} is: {number-2}. ……
…… One of the special magic numbers for {word-3} is: {number-3}. ……
…… One of the special magic numbers for {word-4} is: {number-4}. ……
What is the special magic number for {word-4} mentioned in the provided text?
Task Answer Prefix:
The special magic number for {word-4} mentioned in the provided text is
@@
Task Template:
Some special magic numbers are hidden within the following text. Make sure to memorize it. I will quiz you about the numbers afterwards.
One of the special magic numbers for {word-1} is: {number-1}.
One of the special magic numbers for {word-2} is: {number-2}.
…… One of the special magic numbers for {word-x} is: {number-x}. ……
One of the special magic numbers for {word-n-1} is: {number-n-1}.
One of the special magic numbers for {word-n} is: {number-n}.
What is the special magic number for {word-x} mentioned in the provided text?
Task Answer Prefix:
The special magic number for {word-x} mentioned in the provided text is
@@
Task Template:
Some special magic uuids are hidden within the following text. Make sure to memorize it. I will quiz you about the uuids afterwards.
One of the special magic uuids for {uuid-1} is: {uuid-1}.
One of the special magic uuids for {uuid-2} is: {uuid-2}.
…… One of the special magic uuids for {uuid-x} is: {uuid-x}. ……
One of the special magic uuids for {uuid-n-1} is: {uuid-n-1}.
One of the special magic uuids for {uuid-n} is: {uuid-n}.
What is the special magic number for {uuid-x} mentioned in the provided text?
Task Answer Prefix:
The special magic number for {uuid-x} mentioned in the provided text is
@@
Task Template:
Some special magic numbers are hidden within the following text. Make sure to memorize it. I will quiz you about the numbers afterwards.
Paul Graham Essays.
…… One of the special magic numbers for {word} is: {number-1}. ……
…… One of the special magic numbers for {word} is: {number-2}. ……
…… One of the special magic numbers for {word} is: {number-3}. ……
…… One of the special magic numbers for {word} is: {number-4}. ……
What are all the special magic numbers for {word} mentioned in the provided text?
Task Answer Prefix:
The special magic numbers for {word} mentioned in the provided text are
@@
Task Template:
Some special magic numbers are hidden within the following text. Make sure to memorize it. I will quiz you about the numbers afterwards.
Paul Graham Essays.
…… One of the special magic numbers for {word-1} is: {number-1}. ……
…… One of the special magic numbers for {word-2} is: {number-2}. ……
…… One of the special magic numbers for {word-3} is: {number-3}. ……
…… One of the special magic numbers for {word-4} is: {number-4}. ……
What are all the special magic numbers for {word-1}, {word-2}, {word-3}, and {word-4} mentioned in the provided text?
Task Answer Prefix:
The special magic numbers for {word-1}, {word-2}, {word-3}, and {word-4} mentioned in the provided text are
@@
Task Template:
{one task example}
Memorize and track the chain(s) of variable assignment hidden in the following text.
The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. …… VAR {X1} = {number} ……
…… VAR {X2} = {X1} ……
…… VAR {X3} = {X2} ……
…… VAR {X4} = {X3} ……
…… VAR {X5} = {X4} ……
Question: Find all variables that are assigned the value {number} in the text above.
Task Answer Prefix:
Answer: According to the chain(s) of variable assignment in the text above, 5 variables are assigned the value {number}, they are:
@@
Task Template:
{one task example}
Below is a numbered list of words. In these words, some appear more often than others. Memorize the ones that appear most often.
1. word-a 2. word-b 3. word-c 4. word-a 5. word-d 6. word-a 7. word-e 8. word-f ……
Question: What are the 10 most common words in the above list?
Task Answer Prefix:
Answer: The top 10 words that appear most often in the list are:
@@
Task Template:
Read the following coded text and track the frequency of each coded word. Find the three most frequently appeared coded words. … … word-a … word-b … … … word-c … word-a … word-d word-e … word-a … … word-f … … … … word-g … word-h … word-a … word-i ……
Question: Do not provide any explanation. Please ignore the dots ’….’. What are the three most frequently appeared words in the above coded text?
Task Answer Prefix:
Answer: According to the coded text above, the three most frequently appeared words are:
@@
Task Template:
Answer the question based on the given documents. Only give me the answer and do not output any other words.
The following are given documents.
Document 1:
{document-1}
Document x:
{document-x}
Document n:
{document-n}
Answer the question based on the given documents. Only give me the answer and do not output any other words.
Question: question
Task Answer Prefix:
Answer:
@@
Task Template:
Answer the question based on the given documents. Only give me the answer and do not output any other words.
The following are given documents.
Document 1:
{document-1}
Document x:
{document-x}
Document y:
{document-y}
Document n:
{document-n}
Answer the question based on the given documents. Only give me the answer and do not output any other words.
Question: question
Task Answer Prefix:
Answer:
@@ Appendix E Passkey Retrieval and Vanilla NIAH Results
Command-R (35B) 128k 100.0 100.0 100.0 100.0 100.0 100.0 100.0
@@
Command-R (35B) 128k 100.0 100.0 100.0 100.0 100.0 98.0 99.7

### v3

@@ 7 Conclusion
We present Ruler, a synthetic benchmark for evaluating long-context language models. Ruler contains diverse task categories, retrieval, multi-hop tracing, aggregation and question answering, providing a flexible and comprehensive evaluation of LLM’s long-context capabilities. We benchmark 17 long-context LMs using Ruler with context sizes ranging from 4K to 128K. Despite achieving perfect results in the widely used needle-in-a-haystack test, almost all models fail to maintain their performance in other tasks of Ruler as we increase input length. We observe common failure modes at large context sizes, including the failure to ignore distractors and ineffective utilization of long context (e.g., simply copy from context or use parametric knowledge instead). We show that Ruler is challenging for even the top-ranked open-source models as we increase task complexity. Our analysis further reveals the large potential for improvement on Ruler and the benefit of scaling model sizes in achieving better long context capabilities.
8 Limitations
Despite covering more task categories than retrieval-oriented benchmarks, Ruler is limited in multiple ways which we describe in detail below.
Lack of position controlling.
Current Ruler reports a single number metric for each input length without providing the depth-level performance. The depth-level performance was evaluated by the NIAH test (Kamradt, 2023) and recent works such as LV-Eval (Yuan et al., 2024) and can be effective in revealing the lost-in-the-middle (Liu et al., 2024d) phenomenon. We are aware of this issue and plan to support the position controlling of the key information in our codebase.
Lack of correlation with realistic long-context tasks.
While tasks such as variable tracking and frequent words extraction were proposed to serve as proxies for real long-context natural language tasks, the lack of easy-to-evaluate realistic long-context tasks prevents us from verifying the validity of these proxies. Due to this limitation, we emphasize that Ruler can be used as convenient behavioral checks of long-context language models, however it should not be preferred over more realistic settings, such as NoCHA (Karpinska et al., 2024), which also emphasize on other capabilities such as reasoning and instruction-following.
Lack of evaluation on short context.
In the current Ruler task suite, we include tasks that most models perform reasonably well at 4k context size, and aim to observe performance degradation with the scaling of context size. This should not be misread as perfect LM capabilities at 4k context size. In fact, recent works, such as FlenQA (Levy et al., 2024), have demonstrated degrading performance when increasing their task input length to just a few thousand tokens. While increasing the task complexity in RULER leads to much worse performance at shorter context size, we did not include these results in this paper.
Lack of verification of prompt robustness.
Language models can be sensitive to the prompt format, however we did not extend a comprehensive study on the prompt robustness beyond preliminary testing in the early stage of this work. We also did not heavily experiment with a few fixed hyperparameters in the existing tasks, such as the length of variable names in variable tracking and the synthetic vocabulary size in common word extraction and frequent word extraction.
@@ Appendix A Models
We select in total 37 models for evaluation and analysis. Our results in the main text only include aligned models (GPT-4, Gemini-1.5, and 15 open-source models). Besides the aligned models, we also evaluate 7 open-source base models using Ruler. We use the performance of Llama2-7b (base) and Llama2-7b (chat) at context length of 4K as the threshold for determining effective context size. In our analysis section, we evaluate in total 11 models, including model series Yi and LWM, as well as models of novel architectures, including Mamba and RWKV.
@@
Gemini-1.5 (Reid et al., 2024) - 1M gemini-1.5-pro
Llama3.1 (Meta.AI, 2024b) 70B 128K meta-llama/Meta-Llama-3.1-70B-Instruct
Llama3.1 (Meta.AI, 2024b) 8B 128K meta-llama/Meta-Llama-3.1-8B-Instruct
Command-R-plus (Cohere, 2024) 104B 128K CohereForAI/c4ai-command-r-plus
Qwen2 (Yang et al., 2024) 72B 128K Qwen/Qwen2-72B-Instruct
@@
Mixtral-8x22B (Jiang et al., 2024) 39B/141B 32K mistralai/Mixtral-8x22B-Instruct-v0.1
Mistral-v0.2 (Mistral.AI, 2023) 7B 32K mistralai/Mistral-7B-Instruct-v0.2
GLM4 (GLM et al., 2024) 9B 1M THUDM/glm-4-9b-chat-1m
GradientAI/Llama3 (Meta.AI, 2024a) 70B 1M gradientai/Llama-3-70B-Instruct-Gradient-1048k
Phi3-medium (Abdin et al., 2024) 14B 128K microsoft/Phi-3-medium-128k-instruct
@@
Yi series (Young et al., 2024) 6B,9B 200K 01-ai/Yi-(6B,9B)-200K
LWM series (Liu et al., 2024a) 7B 128K,256K,512K LargeWorldModel/LWM-Text-Chat-(128K,256K,512K)
LWM-base series (Liu et al., 2024a) 7B 32K,128K,256K,512K LargeWorldModel/LWM-Text-(32K,128K,256K,512K)
@@ Appendix B Task Configurations
  • Retrieval: In S-NIAH, we include the passkey retrieval (Mohtashami & Jaggi, 2023) and the vanilla NIAH (Kamradt, 2023), both use word-number as key-value and differ only by the background haystack. Additionally, we change the value type to UUID, for the purpose of testing model robustness at retrieving long strings from context. For MK-NIAH, we add three distractor needles into the haystack. We also include existing setups from previous works: line retrieval (Li et al., 2023a) and key-value retrieval (Liu et al., 2024d) with the haystack filled entirely with distractor needles. For MV-NIAH and MQ-NIAH, we test 4 values and queries respectively.
@@
QA dataset = SQuAD dataset = HotpotQA
@@ Appendix C Task Correlation Analysis
Ruler is designed under the assumption that tasks across different categories are able to reveal distinct model behaviors. We conduct a preliminary correlational study to confirm the validity of task categories and guide the selection of representative tasks. We evaluate eight open-sourced models at various context sizes across 18 task configurations. Each task can then be represented with a vector of model performance at various context sizes. The 18 task vectors are then clustered via agglomorative clustering algorithm, using correlation coefficient as the distance metric. As shown in Figure 5, while certain tasks exhibit moderate correlations with others, tasks in each of the four categories (NIAH, VT, AG, QA) form cohesive clusters of their own without redundancy. We further eliminate a few tasks that correlate highly with other tasks within the same cluster, and finalize 13 tasks for later large scale evaluation.
@@ Appendix D Prompt Templates
GPT-4 {task_template} Do not provide any explanation. Please directly give me the answer. {task_answer_prefix}
Yi/Base {task_template} {task_answer_prefix}
@@
LWM/LongChat {system_prompt} USER: {task_template} ASSISTANT: {task_answer_prefix}
@@
[…]
Llama2/Others [ INST]  {task_template} [ /INST]  {task_answer_prefix}
@@
Task Template: Some special magic numbers are hidden within the following text. Make sure to memorize it. I will quiz you about the numbers afterwards. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. …… One of the special magic numbers for {word} is: {number}. …… What is the special magic number for {word} mentioned in the provided text? Task Answer Prefix: The special magic number for {word} mentioned in the provided text is
@@
Task Template: Some special magic numbers are hidden within the following text. Make sure to memorize it. I will quiz you about the numbers afterwards. Paul Graham Essays. …… One of the special magic numbers for {word} is: {number}. …… What is the special magic number for {word} mentioned in the provided text? Task Answer Prefix: The special magic number for {word} mentioned in the provided text is
@@
Task Template: Some special magic words are hidden within the following text. Make sure to memorize it. I will quiz you about the words afterwards. Paul Graham Essays. …… One of the special magic words for {word} is: {word}. …… What is the special magic word for {word} mentioned in the provided text? Task Answer Prefix: The special magic word for {word} mentioned in the provided text is
@@
Task Template: Some special magic numbers are hidden within the following text. Make sure to memorize it. I will quiz you about the numbers afterwards. Paul Graham Essays. …… One of the special magic numbers for {word-1} is: {number-1}. …… …… One of the special magic numbers for {word-2} is: {number-2}. …… …… One of the special magic numbers for {word-3} is: {number-3}. …… …… One of the special magic numbers for {word-4} is: {number-4}. …… What is the special magic number for {word-4} mentioned in the provided text? Task Answer Prefix: The special magic number for {word-4} mentioned in the provided text is
@@
Task Template: Some special magic numbers are hidden within the following text. Make sure to memorize it. I will quiz you about the numbers afterwards. One of the special magic numbers for {word-1} is: {number-1}. One of the special magic numbers for {word-2} is: {number-2}. …… One of the special magic numbers for {word-x} is: {number-x}. …… One of the special magic numbers for {word-n-1} is: {number-n-1}. One of the special magic numbers for {word-n} is: {number-n}. What is the special magic number for {word-x} mentioned in the provided text? Task Answer Prefix: The special magic number for {word-x} mentioned in the provided text is
@@
Task Template: Some special magic uuids are hidden within the following text. Make sure to memorize it. I will quiz you about the uuids afterwards. One of the special magic uuids for {uuid-1} is: {uuid-1}. One of the special magic uuids for {uuid-2} is: {uuid-2}. …… One of the special magic uuids for {uuid-x} is: {uuid-x}. …… One of the special magic uuids for {uuid-n-1} is: {uuid-n-1}. One of the special magic uuids for {uuid-n} is: {uuid-n}. What is the special magic number for {uuid-x} mentioned in the provided text? Task Answer Prefix: The special magic number for {uuid-x} mentioned in the provided text is
@@
Task Template: Some special magic numbers are hidden within the following text. Make sure to memorize it. I will quiz you about the numbers afterwards. Paul Graham Essays. …… One of the special magic numbers for {word} is: {number-1}. …… …… One of the special magic numbers for {word} is: {number-2}. …… …… One of the special magic numbers for {word} is: {number-3}. …… …… One of the special magic numbers for {word} is: {number-4}. …… What are all the special magic numbers for {word} mentioned in the provided text? Task Answer Prefix: The special magic numbers for {word} mentioned in the provided text are
@@
Task Template: Some special magic numbers are hidden within the following text. Make sure to memorize it. I will quiz you about the numbers afterwards. Paul Graham Essays. …… One of the special magic numbers for {word-1} is: {number-1}. …… …… One of the special magic numbers for {word-2} is: {number-2}. …… …… One of the special magic numbers for {word-3} is: {number-3}. …… …… One of the special magic numbers for {word-4} is: {number-4}. …… What are all the special magic numbers for {word-1}, {word-2}, {word-3}, and {word-4} mentioned in the provided text? Task Answer Prefix: The special magic numbers for {word-1}, {word-2}, {word-3}, and {word-4} mentioned in the provided text are
@@
Task Template: {one task example} Memorize and track the chain(s) of variable assignment hidden in the following text. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. …… VAR {X1} = {number} …… …… VAR {X2} = {X1} …… …… VAR {X3} = {X2} …… …… VAR {X4} = {X3} …… …… VAR {X5} = {X4} …… Question: Find all variables that are assigned the value {number} in the text above. Task Answer Prefix: Answer: According to the chain(s) of variable assignment in the text above, 5 variables are assigned the value {number}, they are:
@@
Task Template: {one task example} Below is a numbered list of words. In these words, some appear more often than others. Memorize the ones that appear most often. 1. word-a 2. word-b 3. word-c 4. word-a 5. word-d 6. word-a 7. word-e 8. word-f …… Question: What are the 10 most common words in the above list? Task Answer Prefix: Answer: The top 10 words that appear most often in the list are:
@@
Task Template: Read the following coded text and track the frequency of each coded word. Find the three most frequently appeared coded words. … … word-a … word-b … … … word-c … word-a … word-d word-e … word-a … … word-f … … … … word-g … word-h … word-a … word-i …… Question: Do not provide any explanation. Please ignore the dots ’….’. What are the three most frequently appeared words in the above coded text? Task Answer Prefix: Answer: According to the coded text above, the three most frequently appeared words are:
@@
Task Template: Answer the question based on the given documents. Only give me the answer and do not output any other words. The following are given documents. Document 1: {document-1} …… Document x: {document-x} …… Document n: {document-n} Answer the question based on the given documents. Only give me the answer and do not output any other words. Question: question Task Answer Prefix: Answer:
@@
Task Template: Answer the question based on the given documents. Only give me the answer and do not output any other words. The following are given documents. Document 1: {document-1} …… Document x: {document-x} …… Document y: {document-y} …… Document n: {document-n} Answer the question based on the given documents. Only give me the answer and do not output any other words. Question: question Task Answer Prefix: Answer:
@@ Appendix E Passkey Retrieval and Vanilla NIAH Results
Llama3.1 (70B) 128K 100.0 100.0 100.0 100.0 100.0 97.8 99.6
Llama3.1 (8B) 128K 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Qwen2 (72B) 128K 100.0 100.0 100.0 100.0 100.0 96.6 99.4
Command-R-plus (104B) 128K 100.0 100.0 99.8 99.8 100.0 97.2 99.5
GLM4 (9B) 1M 100.0 100.0 100.0 100.0 100.0 100.0 100.0
GradientAI/Llama3 (70B) 1M 100.0 100.0 100.0 100.0 100.0 93.6 98.9
Mixtral-8x22B (39B/141B) 64K 100.0 100.0 100.0 100.0 99.6 0.0 83.3
@@
Llama3.1 (70B) 128K 100.0 100.0 100.0 100.0 100.0 99.6 99.9
Llama3.1 (8B) 128K 100.0 100.0 100.0 100.0 100.0 99.6 99.9
Qwen2 (72B) 128K 100.0 100.0 100.0 100.0 99.8 56.4 92.7
Command-R-plus (35B) 128K 100.0 100.0 100.0 100.0 99.8 86.0 97.6
GLM4 (9B) 128K 100.0 100.0 100.0 100.0 100.0 100.0 100.0
GradientAI/Llama3 (70B) 1M 100.0 100.0 100.0 99.6 99.2 97.8 99.4
Mixtral-8x22B (39B/141B) 64K 100.0 100.0 100.0 100.0 99.6 24.2 87.3